The radius of cylinder A is 4m, Cylinder B is 6m. Solve for reaction A and B
1
Expert's answer
2021-08-26T08:10:57-0400

Volume of cylinder A = "\dfrac{4}{3}\pi r^3=268.08\space m^3"

Volume of cylinder B = "\dfrac{4}{3}\pi r^3=904.77\space m^3"